# Sampling by averaging: A multiscale approach to score estimation

### v1

@@ 3.2 MultCDiff: Multiscale Controlled Diffusions
Annealed Langevin dynamics (3) offers a convenient approach to sampling from the target distribution for general schedules $\lambda_{t}$ that interpolate between the base and target distributions in finite time. However, a key drawback of annealed Langevin dynamics is that it introduces bias—even if perfectly simulated (see Section 4.2). This bias can be corrected by incorporating control terms. Although such control terms are generally intractable for arbitrary schedules, they can be computed explicitly in the case of an OU noising process. The multiscale approach introduced in the previous section naturally extends to this controlled setting, in both overdamped and underdamped regimes. Given the improved efficiency of underdamped dynamics, we focus our discussion on the underdamped formulation (see App. B.2.1 for details on the overdamped case). When using an OU process as the forward noising model, the underdamped diffusion, initialised at $\overrightarrow{X}_{0}\sim\pi$ and $\overrightarrow{V}_{0}\sim\mathcal{N}(0,I)$, is defined as
@@
In this case, to ensure efficient and stable dynamics, we adopt the critical damping regime, where $\Gamma^{2}=4M$. This choice leads to fast convergence without oscillations [McCall, 2010]. The corresponding time-reversed SDE is
$\begin{cases}\mathrm{d}\overleftarrow{X}_{t}=-4\Gamma^{-2}\overleftarrow{V}_{t}\ \mathrm{d}t\\ \mathrm{d}\overleftarrow{V}_{t}=\left(\overleftarrow{X}_{t}+4\Gamma^{-1}\overleftarrow{V}_{t}+2\Gamma\nabla_{V_{t}}\log q_{T-t}(\overleftarrow{X}_{t},\overleftarrow{V}_{t})\right)\mathrm{d}t+\sqrt{2\Gamma}\ \mathrm{d}B_{t},\end{cases}$ (8)
@@
To implement this sampler—referred to as MultCDiff—we construct a novel integrator inspired by Dockhorn et al. [2022b]. Specifically, we leverage the symmetric Trotter splitting and the Baker–Campbell–Hausdorff formula [Trotter, 1959; Strang, 1968; Tuckerman, 2010] to design a stable and efficient symmetric splitting scheme, see App. B.2 for details. We also outline in the appendix the difficulties of extending controlled diffusions to the heavy-tailed scenario.
@@ 4 Theoretical guarantees
In this section, we identify the different sources of error in the proposed sampling algorithms. For MultALMC, the total error consists of three components: the discretisation error from numerically solving the slow-fast system, the convergence error of the slow-fast system to its averaged dynamics, and the bias arising from the time-inhomogeneous averaged system, whose marginals differ from those of the diffusion path in (1)—this last component is discussed in 4.2. For MultCDiff, the error bound includes an initialisation error present in the error bounds of diffusion models [Benton et al., 2024; Chen et al., 2023], the discretisation error of the slow-fast system, and the convergence of the multiscale system to the averaged dynamics. The discretisation error depends on the choice of numerical integrator, underscoring the importance of an accurate and stable integrator scheme. We do not analyse this error in detail as it follows established methods [Monmarché, 2020; Leimkuhler et al., 2024]. We now study the convergence properties of the slow component $X_{t}$ to the solution of the corresponding averaged dynamics that evolves along the time-inhomogeneous diffusion path in (1).
@@ 4.1 Convergence of the slow-fast system to the averaged dynamics
Building on results from stochastic averaging theory [Liu et al., 2020], our goal is to derive sufficient conditions on the coefficients of the multiscale systems—namely, (4), (6), and (10)—that guarantee strong convergence of the slow component to the averaged process as $\varepsilon\to 0$. We focus on the case where the base distribution is Gaussian $\nu\sim\mathcal{N}(0,I)$. Extending the convergence analysis to the case where $\nu$ is a heavy-tailed distribution, such as a Student’s t distribution, presents significant technical challenges, which fall outside of the scope of this work. We outline these challenges in App. C.2. To establish convergence guarantees, we assume the following regularity conditions.
@@
The schedule $\lambda_{t}:\mathbb{R}^{+}\to[0,1]$ is a non-decreasing function of $t$, weakly differentiable and Hölder continuous with constant $\gamma_{1}$ in $(0,1]$.
@@ 4.2 Bias of the annealed Langevin dynamics
The bias of the overdamped annealed Langevin dynamics (3) has been studied in prior work [Guo et al., 2025; Cordero-Encinar et al., 2025]. Here, we focus on quantifying the bias introduced by the underdamped averaged system (5) relative to the true diffusion path in the augmented state space with the velocity. At fixed time $t$, the invariant distribution of the underdamped averaged system takes the form
@@
Note that the Hamiltonian is separable, meaning that $x$ and $v$ are independent. It is important to emphasise that, even when simulated exactly, diffusion annealed Langevin dynamics introduces a bias, as the marginal distributions of the solution of (5) do not exactly match the prescribed distributions $(\hat{p}_{t})_{t}$. A key quantity for characterising this bias will be the action of the curve of probability measures $\mu=(\mu_{t})_{t\in[0,1]}$ defined in (1) interpolating between the base distribution $\nu$ and the target distribution $\pi$, denoted by $\mathcal{A}(\mu)$. As noted by Guo et al. [2025], the action serves as a measure of the cost of transporting $\nu$ to $\pi$ along the specified path. Formally, for an absolutely continuous curve of probability measures [Lisini, 2007] with finite second-order moment, the action is given by
$\mathcal{A}(\mu):=\int_{0}^{T}\lim_{\delta\to 0}{|\delta|}^{-1}{W_{2}(\mu_{t+\delta},\mu_{t})}\mathrm{d}{t}.$
Based on Theorem 1 from Guo et al. [2025], we have the following result.
@@
The action $\mathcal{A}(\mu)$ is bounded when the target $\pi$ has bounded second order moment and under mild conditions on the schedule function, see Cordero-Encinar et al. [2025, Lemmas 3.3 and 4.2]. Full details and a complete proof are provided in App. C.3.
@@ 5 Numerical experiments
  • Funnel: 10-dimensional distribution defined by $\pi(x)\propto\mathcal{N}(x_{1};0,\eta^{2})\prod_{i=2}^{d}\mathcal{N}(x_{i};0,e^{x_{1}})$ for $x=(x_{i})_{i=1}^{10}\in\mathbb{R}^{10}$ with $\eta=3$ [Neal, 2003].
  • Double well potential (DW): $d$-dimensional distribution given by $\pi\propto\exp(-\sum_{i=1}^{m}(x_{i}^{2}-\delta)^{2}-\sum_{i=m+1}^{d}x_{i}^{2})$ with $m\in\mathbb{N}$ and $\delta\in(0,\infty)$. Ground truth samples are obtained using rejection sampling with a Gaussian mixture proposal distribution [Midgley et al., 2023].
@@
  • Statistical physics model: Sampling metastable states of the stochastic Allen-Cahn equation $\phi^{4}$ model (dimension 100) [Albergo et al., 2019; Gabrié et al., 2022]. At the chosen temperature, the distribution has two well distinct modes with relative weight controlled by a parameter $h$. Following Grenioux et al. [2024], we estimate the relative weight between the two modes.
@@
We compare our approach against a representative selection of related sampling methods: Sequential Monte Carlo (SMC) [Del Moral et al., 2006; Doucet et al., 2001], Annealed Importance Sampling (AIS) [Neal, 2001], Underdamped Langevin Monte Carlo (ULMC) [Cheng et al., 2018; Neal et al., 2011], Parallel Tempering (PT) [Swendsen and Wang, 1986; Geyer and Thompson, 1995], Diffusive Gibbs Sampling (DiGS) [Chen et al., 2024], Reverse Diffusion Monte Carlo (RDMC) [Huang et al., 2024] and Stochastic Localisation via Iterative Posterior Sampling (SLIPS) [Grenioux et al., 2024]. We ensure a fair comparison by using the same number of energy evaluations and the same initialisation based on a standard Gaussian distribution for all baselines. For completeness, we also report the performance of SLIPS with optimal initialisation (App. D)—which assumes knowledge of the distribution’s scalar variance, unavailable in practice. For the $\phi^{4}$ model, we only compare our results against a Laplace approximation and SLIPS algorithm, as the other methods provide degenerate samples. Since our approach is learning-free, we exclude comparisons with neural-based samplers as it would be difficult to make comparisons at equal computational budget.
@@ Appendix A Preliminaries
Before presenting a formal definition, we first provide the intuition behind multiscale methods, following Schuh and Souttar [2024]. In many areas of science and engineering [Bertram and Rubin, 2017; Weinan et al., 2005], one often encounters systems that are difficult to analyse directly. However, there may exist a related system that is more tractable, though it does not produce exactly the same behaviour. In such cases, the simpler system can be viewed as an approximation of the more complex one. In our setting, the simpler system is a slow-fast system, while the complex system corresponds to the averaged dynamics. We develop these concepts in the following.
@@
$\begin{cases}\mathrm{d}X_{t}^{\varepsilon}=b(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}t+\sigma(t,X_{t}^{\varepsilon})\mathrm{d}B_{t}\quad&X_{0}^{\varepsilon}=x\in\mathbb{R}^{n}\\ \mathrm{d}Y_{t}^{\varepsilon}=\frac{1}{\varepsilon}f(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}t+\frac{1}{\sqrt{\varepsilon}}g(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}\tilde{B}_{t}\quad&Y_{0}^{\varepsilon}=y\in\mathbb{R}^{m},\end{cases}$ (12)
@@
$\alpha_{T}(R):=\int_{0}^{T}K_{t}(R)\mathrm{d}t<\infty,\;\;\text{on }\Omega,\quad\mathbb{E}\left[e^{p\alpha_{T}(1)}\right]<\infty,\quad\sup_{t\in[0,T]}\mathbb{E}\|K_{t}(1)\|^{4}<\infty.$
@@ B.1.1 Overdamped system
Note that if the slow-fast system converges to its corresponding averaged dynamics, then running the modified multiscale system up to time $t_{\delta}$, where $\lambda_{\kappa t_{\delta}}=1-\lambda_{\delta}$, we can approximate the distribution $\hat{\mu}_{t_{\delta}}$ that is close to the target $\pi$, provided $\lambda_{\delta}$ is sufficiently small. For $t\geq t_{\delta}$, the dynamics in (13) reduce to standard overdamped Langevin dynamics targetting $\pi$, initialised at $\hat{\mu}_{t_{\delta}}$. This warm start is known to significantly improve convergence rates to the target distribution [Wu et al., 2022; Lee et al., 2021a; Chewi et al., 2021].
@@
$\quad\quad\quad\quad X_{l+1}=\begin{cases}X_{l}+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l}-Y_{l}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)h_{l}+\sqrt{2h_{l}}\xi_{l}^{(1)}&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ X_{l}-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)h_{l}+\sqrt{2h_{l}}\xi_{l}^{(1)}&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
@@
$\quad\quad\quad\quad\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,0}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,0}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,i-1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,i-1}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l}^{(2)}\end{cases}$

### v2

@@ 3.2 MultCDiff: Multiscale Controlled Diffusions
Annealed Langevin dynamics (3) offers a convenient approach to sampling from the target distribution for general schedules $\lambda_{t}$ that interpolate between the base and target distributions in finite time. However, a key drawback of annealed Langevin dynamics is that it introduces bias, even if perfectly simulated (see Section 4.2). This bias can be corrected by incorporating control terms. Although such control terms are generally intractable for arbitrary schedules, they can be computed explicitly in the case of an OU noising process. The multiscale approach introduced in the previous section naturally extends to this controlled setting, in both overdamped and underdamped regimes. Given the improved efficiency of underdamped dynamics, we focus our discussion on the underdamped formulation (see App. B.2.1 for details on the overdamped case). When using an OU process as the forward noising model, the underdamped diffusion, initialised at $\overrightarrow{X}_{0}\sim\pi$ and $\overrightarrow{V}_{0}\sim\mathcal{N}(0,I)$, is defined as
@@
In this case, to ensure efficient and stable dynamics, we adopt the critical damping regime, where $\Gamma^{2}=4M$. This choice leads to fast convergence without oscillations (McCall, 2010). The corresponding time-reversed SDE is
$\begin{cases}\mathrm{d}\overleftarrow{X}_{t}=-4\Gamma^{-2}\overleftarrow{V}_{t}\ \mathrm{d}t\\ \mathrm{d}\overleftarrow{V}_{t}=\left(\overleftarrow{X}_{t}+4\Gamma^{-1}\overleftarrow{V}_{t}+2\Gamma\nabla_{v}\log q_{T-t}(\overleftarrow{X}_{t},\overleftarrow{V}_{t})\right)\mathrm{d}t+\sqrt{2\Gamma}\ \mathrm{d}B_{t},\end{cases}$ (8)
@@
To implement this sampler, referred to as MultCDiff, we construct a novel integrator inspired by Dockhorn et al. (2022b). Specifically, we leverage the symmetric Trotter splitting and the Baker–Campbell–Hausdorff formula (Trotter, 1959; Strang, 1968; Tuckerman, 2010) to design a stable and efficient symmetric splitting scheme, see App. B.2 for details. We also outline in the appendix the difficulties of extending controlled diffusions to the heavy-tailed scenario.
@@ 4 Theoretical guarantees
In this section, we identify the different sources of error in the proposed sampling algorithms. For MultALMC, the total error consists of three components: the discretisation error from numerically solving the slow-fast system, the convergence error of the slow-fast system to its averaged dynamics, and the bias arising from the time-inhomogeneous averaged system, whose marginals differ from those of the diffusion path in (1)—this last component is discussed in 4.2. For MultCDiff, the error bound includes an initialisation error present in the error bounds of diffusion models (Benton et al., 2024; Chen et al., 2023), the discretisation error of the slow-fast system, and the convergence of the multiscale system to the averaged dynamics. The discretisation error depends on the choice of numerical integrator, underscoring the importance of an accurate and stable integrator scheme. We do not analyse this error in detail as it follows established methods (Monmarché, 2020; Leimkuhler et al., 2024). We now study the convergence properties of the slow component $X_{t}$ to the solution of the corresponding averaged dynamics that evolves along the time-inhomogeneous diffusion path in (1).
@@ 4.1 Convergence of the slow-fast system to the averaged dynamics
Building on results from stochastic averaging theory (Liu et al., 2020), our goal is to derive sufficient conditions on the coefficients of the multiscale systems—namely, (4), (6), and (10)—that guarantee strong convergence of the slow component to the averaged process as $\varepsilon\to 0$. We focus on the case where the base distribution is Gaussian $\nu\sim\mathcal{N}(0,I)$. Extending the convergence analysis to the case where $\nu$ is a heavy-tailed distribution, such as a Student’s t distribution, presents significant technical challenges, which fall outside of the scope of this work. We outline these challenges in App. C.2. To establish convergence guarantees, we assume the following regularity conditions.
@@
The schedule $\lambda_{t}:\mathbb{R}^{+}\to[0,1]$ is a non-decreasing function of $t$, weakly differentiable and Hölder continuous with exponent $\gamma_{1}$ in $(0,1]$.
@@ 4.2 Bias of the annealed Langevin dynamics
The bias of the overdamped annealed Langevin dynamics (3) has been studied in prior work (Guo et al., 2025; Cordero-Encinar et al., 2025). Here, we focus on quantifying the bias introduced by the underdamped averaged system (5) relative to the true diffusion path in the augmented state space with the velocity. At fixed time $t$, the invariant distribution of the underdamped averaged system takes the form
@@
Note that the Hamiltonian is separable, meaning that $x$ and $v$ are independent. It is important to emphasise that, even when simulated exactly, diffusion annealed Langevin dynamics introduces a bias, as the marginal distributions of the solution of (5) do not exactly match the prescribed distributions $(\hat{p}_{t})_{t}$. A key quantity for characterising this bias will be the action of the curve of probability measures $\mu=(\mu_{t})_{t\in[0,1]}$ defined in (1) interpolating between the base distribution $\nu$ and the target distribution $\pi$, denoted by $\mathcal{A}(\mu)$. As noted by Guo et al. (2025), the action serves as a measure of the cost of transporting $\nu$ to $\pi$ along the specified path. Formally, for an absolutely continuous curve of probability measures (Lisini, 2007) with finite second-order moment, the action is given by
$\mathcal{A}(\mu):=\int_{0}^{T}\left(\lim_{\delta\to 0}{|\delta|}^{-1}{W_{2}(\mu_{t+\delta},\mu_{t})}\right)^{2}\mathrm{d}{t}.$
Based on Theorem 1 from Guo et al. (2025), we have the following result.
@@
The action $\mathcal{A}(\mu)$ is bounded when the target $\pi$ has bounded second order moment and under mild conditions on the schedule function, see Cordero-Encinar et al. (2025, Lemmas 3.3 and 4.2). Full details and a complete proof are provided in App. C.3.
@@ 5 Numerical experiments
  • Funnel: 10-dimensional distribution defined by $\pi(x)\propto\mathcal{N}(x_{1};0,\eta^{2})\prod_{i=2}^{d}\mathcal{N}(x_{i};0,e^{x_{1}})$ for $x=(x_{i})_{i=1}^{10}\in\mathbb{R}^{10}$ with $\eta=3$ (Neal, 2003).
  • Double well potential (DW): $d$-dimensional distribution given by $\pi\propto\exp(-\sum_{i=1}^{m}(x_{i}^{2}-\delta)^{2}-\sum_{i=m+1}^{d}x_{i}^{2})$ with $m\in\mathbb{N}$ and $\delta\in(0,\infty)$. Ground truth samples are obtained using rejection sampling with a Gaussian mixture proposal distribution (Midgley et al., 2023).
@@
  • Statistical physics model: Sampling metastable states of the stochastic Allen-Cahn equation $\phi^{4}$ model (dimension 100) (Albergo et al., 2019; Gabrié et al., 2022). At the chosen temperature, the distribution has two well distinct modes with relative weight controlled by a parameter $h$. Following Grenioux et al. (2024), we estimate the relative weight between the two modes.
@@
We compare our approach against a representative selection of related sampling methods: Sequential Monte Carlo (SMC) (Del Moral et al., 2006; Doucet et al., 2001), Annealed Importance Sampling (AIS) (Neal, 2001), Underdamped Langevin Monte Carlo (ULMC) (Cheng et al., 2018; Neal et al., 2011), Parallel Tempering (PT) (Swendsen and Wang, 1986; Geyer and Thompson, 1995), Diffusive Gibbs Sampling (DiGS) (Chen et al., 2024), Reverse Diffusion Monte Carlo (RDMC) (Huang et al., 2024) and Stochastic Localisation via Iterative Posterior Sampling (SLIPS) (Grenioux et al., 2024). We ensure a fair comparison by using the same number of energy evaluations and the same initialisation based on a standard Gaussian distribution for all baselines. For completeness, we also report the performance of SLIPS with optimal initialisation (App. D)—which assumes knowledge of the distribution’s scalar variance, unavailable in practice. For the $\phi^{4}$ model, we only compare our results against a Laplace approximation and SLIPS algorithm, as the other methods provide degenerate samples. Since our approach is learning-free, we exclude comparisons with neural-based samplers as it would be difficult to make comparisons at equal computational budget.
@@ Appendix A Preliminaries
Before presenting a formal definition, we first provide the intuition behind multiscale methods, following Schuh and Souttar [2024]. In many areas of science and engineering [Bertram and Rubin, 2017, Weinan et al., 2005], one often encounters systems that are difficult to analyse directly. However, there may exist a related system that is more tractable, though it does not produce exactly the same behaviour. In such cases, the simpler system can be viewed as an approximation of the more complex one. In our setting, the simpler system is a slow-fast system, while the complex system corresponds to the averaged dynamics. We develop these concepts in the following.
@@
$\begin{cases}\mathrm{d}X_{t}^{\varepsilon}=b(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}t+\sigma(t,X_{t}^{\varepsilon})\mathrm{d}B_{t}\qquad&X_{0}^{\varepsilon}=x\in\mathbb{R}^{n}\\ \mathrm{d}Y_{t}^{\varepsilon}=\frac{1}{\varepsilon}f(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}t+\frac{1}{\sqrt{\varepsilon}}g(t,X_{t}^{\varepsilon},Y_{t}^{\varepsilon})\mathrm{d}\tilde{B}_{t}\qquad&Y_{0}^{\varepsilon}=y\in\mathbb{R}^{m},\end{cases}$ (12)
@@
$\alpha_{T}(R):=\int_{0}^{T}K_{t}(R)\mathrm{d}t<\infty,\;\;\text{on }\Omega,\qquad\mathbb{E}\left[e^{p\alpha_{T}(1)}\right]<\infty,\qquad\sup_{t\in[0,T]}\mathbb{E}\|K_{t}(1)\|^{4}<\infty.$
@@ B.1.1 Overdamped system
Note that if the slow-fast system converges to its corresponding averaged dynamics, then running the modified multiscale system up to time $t_{\delta}$, where $\lambda_{\kappa t_{\delta}}=1-\lambda_{\delta}$, we can approximate the distribution $\hat{\mu}_{t_{\delta}}$ that is close to the target $\pi$, provided $\lambda_{\delta}$ is sufficiently small. For $t\geq t_{\delta}$, the dynamics in (13) reduce to standard overdamped Langevin dynamics targetting $\pi$, initialised at $\hat{\mu}_{t_{\delta}}$. This warm start is known to significantly improve convergence rates to the target distribution [Wu et al., 2022, Lee et al., 2021a, Chewi et al., 2021].
@@
$\qquad\qquad\qquad\qquad X_{l+1}=\begin{cases}X_{l}+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l}-Y_{l}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)h_{l}+\sqrt{2h_{l}}\xi_{l}^{(1)}&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ X_{l}-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)h_{l}+\sqrt{2h_{l}}\xi_{l}^{(1)}&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
@@
$\qquad\qquad\qquad\qquad\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,0}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,0}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,i-1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,i-1}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l}^{(2)}\end{cases}$